OpenStudy (anonymous):

What is the slope of the line which passes through (–2, 0) and (0, 4)?

OpenStudy (hoblos):

(4-0)/0+2) = 2

OpenStudy (nenadmatematika):

it's (4-0)/(0+2)=2 :D

OpenStudy (anonymous):

a) w = P + 2l – 2 b) w = P + 2l c) w= 2l-P/2 d) w= P-2l/2

OpenStudy (anonymous):

slope is 2

OpenStudy (anonymous):

0-4=-4 -2-0=-2 slope =-4/-2=2 final ans

OpenStudy (anonymous):

w=p+21-2; w=p+19